It's still improving, now the current is 0.3 mA (0.38mA yesterday). I'll wait until it stops improving and try to tighten the screws a bit. \
About the plastic pieces, they are peek. I think it is vacuum compatible, cf E960050-v11.
May be it is the temp sensor that I have to re-solder on the copper pieces. I did not bake them after soldering.
The gap at the frange is OK as long as the gasket is evenly squeezed.
1. It the pressure still improving? => Your cavity is still out gassing.
2. Fasten the screws little more (not too much) for these two franges.
=> If there is any improvement there is some leakage.
=> If not, this is just a outgas from your cavity.
Do you think your black plastic pieces are vacuum compatible? Are they made of Delrin?